Improper Integral (Integrals with Infinite Limits)


∫dz/(z-1)^(2/3) from 0 to 2


1
Expert's answer
2021-05-07T13:51:01-0400

Given integral is-


"\int_0^2\dfrac{dz}{(z-1)^{\frac{2}{3}}}



\\[9pt]=\dfrac{(z-1)^{\frac{1}{3}}}{\frac{1}{3}}|_0^2



\\[9pt]=3(2-1)^{\frac{1}{3}}-3(0-1)^{\frac{1}{3}}



\\[9pt]=3(1)-3(-1)^{\frac{1}{3}}\\[9pt]=3(1-(-1)^{\frac{1}{3}})"